aqueous sodium chloride can be electrolysed using inert electrodes. Describe the effect that the position of ions in the electrochemical series has on the products of this electrolysis


1
Expert's answer
2021-02-23T04:34:46-0500


The Na+ ions migrate toward the negative electrode and the Cl- ions migrate toward the positive electrode. But, now there are two substances that can be reduced at the cathode: Na+ ions and water molecules.

 Electrolysis:

"2NaCl(l)\rightarrow2Na(l)+Cl_2(g)"



Cathode (-):    

"Na^+ + e^-\rightarrow Na" "E^{\degree}_{red}= -2.71V"

 "2H_2O + 2e^-\rightarrow H_2+2OH^-" "E^{\degree}_{red}=-0.83V"


Because it is much easier to reduce water than Na+ ions, the only product formed at the cathode is hydrogen gas.

So, final product on cathode is:

Cathode (-): "2H_2O(l)+2e^- \rightarrow H_2(g)+2OH^-(aq)"


There are also two substances that can be oxidized at the anode: Cl- ions and water molecules.



Anode (+):   

  "2Cl^-\rightarrow Cl_2 +2e^-" "E^{\degree}_{ox}= -1.36V"  

"2H_2O\rightarrow O_2+4H^++4e^-" "E^{\degree}_{ox}= -1.23V"


The standard-state potentials for these half-reactions are so close to each other that we might expect to see a mixture of Cl2 and O2 gas collect at the anode. In practice, the only product of this reaction is Cl2.

So. final product on anode is:


Anode (+): "2Cl^-\rightarrow Cl_2+ 2e^-"


Electrolysis of aqueous "NaCl" solutions gives a mixture of hydrogen and chlorine gas and an aqueous sodium hydroxide solution.

  Final Electrolysis:

"2NaCl(aq)+2H_2O(l)\rightarrow 2Na^+(aq)+2OH^-(aq)+H_2(g)+Cl_2(g)"